# AFABench: A Generic Framework for Benchmarking Active Feature Acquisition

### v1

@@ 3.1.4. Static feature selection methods
3.2. Datasets
Table 2. Summary of datasets used in the benchmark.
Dataset Type Train Size Val Size Test Size # Features # Classes
CUBE Synthetic 700 150 150 20 8
AFAContext Synthetic 700 150 150 30 8
MNIST Real World 42,000 9,000 9,000 784 10
FashionMNIST Real World 42,000 9,000 9,000 784 10
Diabetes Real World 64,443 13,810 13,810 45 3
PhysioNet Real World 8,400 1,800 1,800 41 2
MiniBooNE Real World 91,044 19,509 19,511 50 2
In this section, we introduce the datasets used in our experiments. A summary of all datasets is provided in Table 2. CUBE is a widely used synthetic dataset in the AFA literature. We describe its structure and limitations in Section 4, and introduce a new synthetic dataset, AFAContext, to address an important shortcoming of CUBE. MNIST is the standard handwritten digit classification dataset (LeCun et al., 1998). Following prior work in AFA (Gadgil et al., 2024), we treat MNIST as a tabular task by considering each pixel as a separate feature. FashionMNIST is a dataset of grayscale images of clothing items such as shirts, trousers, and shoes (Xiao et al., 2017). It shares the same format and image dimensions as MNIST, and we similarly treat each pixel as a separate feature for tabular modeling. Diabetes is a real-world dataset for diabetes diagnosis, making it naturally suitable for AFA since different patients may require different examinations for an accurate diagnosis. This dataset has been widely used in previous AFA research (Kachuee et al., 2019; Covert et al., 2023). As in previous studies, we define three classes—normal, pre-diabetes, and diabetes—based on standard fasting glucose thresholds. The data originates from the National Health and Nutrition Examination Survey (NHANES) (NHA, 2018), a large-scale, ongoing health survey of U.S. adults and children. We use the pre-processed version made available by Kachuee et al. (2019), which includes data collected between 1999 and 2016. PhysioNet is another medical dataset, derived from the PhysioNet Challenge 2012 (Goldberger et al., 2000). It contains electronic health records (EHR) from ICU patients, with the goal of predicting in-hospital mortality, a binary classification task—based on results from various clinical tests and measurements. MiniBooNE is a particle identification dataset collected by the MiniBooNE experiment at Fermilab (Roe, 2005), where the task is to distinguish signal events (electron neutrinos) from background events (muon neutrinos) based on reconstructed particle interaction features. The dataset is fully tabular, each instance is represented by a set of real-valued detector-level features.
3.3. Evaluation Protocol
We report the performance of each method ($y$-axis) as a function of the budget, i.e., number of selected features ($x$-axis). Accuracy is used as the evaluation metric for all datasets, except PhysioNet, where we report F1-score due to its high class imbalance.
3.4. Setup for Fair Benchmarking
In this section, we discuss the assumptions necessary to provide a benchmark that is as fair and objective as possible.
Tabular data. All experiments are conducted on tabular datasets (see Section 3.2 for details), as some of the included methods do not scale well to other domains, such as image data. For example, EDDI-GG is known to not scale well beyond tabular datasets (Gadgil et al., 2024).
Hard budget. We assume a fixed feature budget $b$ for each test instance. Although many of the included methods (see Section 3.1) can be extended to the variable-budget setting—by allowing the policy $\pi$ to return a stop action indicating when to stop acquisition and make a prediction, such extensions complicate a fair comparison. For example, RL-based methods can incorporate a stop action in the action space (Janisch et al., 2020), while greedy methods might stop based on confidence thresholds, such as when the entropy of $p(\mathbf{y}\mid x_{S})$ is sufficiently low. However, these methods require tuning additional and often unintuitive hyperparameters to balance feature cost and predictive performance. Because each method handles this trade-off differently, it becomes difficult to ensure a fair comparison. Setting a fixed hard budget $b$ for all methods allows for a more controlled and objective evaluation. Moreover, specifying a fixed budget is often more practical in real-world scenarios, where users may find it easier to define a maximum number of features rather than fine-tuning cost-performance trade-offs (Janisch et al., 2020).
Common components. All methods are implemented in a unified framework, where common components between the methods are consistent. We now provide some examples of this. (i) Both EDDI-GG and ODIN-MBRL use a pre-trained PVAE (although for different purposes). We use the same pre-trained model for both. (ii) We use a shared pre-trained predictor $f$ for each task. However, some methods are specifically designed to train a predictor $f$ jointly with the policy $\pi$. In such cases, we report the performance of both the shared classifier and the one that was jointly trained.
Consistent RL framework Agents in RL-based methods interact with the same MDP, apart from the reward function, which is method-dependent.
Excluded methods. As discussed in Rahbar et al., there are other approaches to AFA that go beyond those included in our benchmark. These include (i) embedded methods and (ii) methods based on adaptive submodular objectives, such as EC2 (Golovin and Krause, 2011). Embedded methods incorporate feature acquisition directly into the model’s training or inference process—e.g., decision trees, random forests, or Naive Bayes—by modifying splitting criteria or loss functions to account for acquisition cost. However, such methods are tightly coupled to specific model architectures and cannot be employed with the shared predictor $f$ (a fully-connected neural network) used in our experiments. Methods based on adaptive submodularity focus on theoretical guarantees, but often rely on strong assumptions about the data distribution, model class, or cost structure. In contrast, we focus on more general methods that are model-agnostic and compatible with a wide range of domains and predictors.
Randomness. In our experiments, we report the average performance over (i) multiple data splits (i.e., training, validation, and test), and (ii) multiple runs with different random seeds for retraining the policy $\pi$, ensuring consistent and unified evaluation across all methods. We also include the standard deviation across runs, visualized as error bars.
Adherence to original architecture Although the most fair approach would be to perform extensive hyperparameter sweeps for each combination of methods/data sets/budgets, this is not realistic in practice. Therefore, we have opted to use the hyperparameters presented in the original paper / repository for each method, although some exceptions to this rule do exist. For example, when the parameters of a method have been tuned to datasets with a small number of features and we find that it underfits to (Fashion)MNIST, we have increased the complexity of the model. Some other changes have been made for consistency reasons. A list of the most important changes made compared to the original implementation can be found in Appendix A.
(a) CUBE
(b) AFAContext
Figure 1. Visualization of (a) CUBE, and (b) our synthetic dataset AFAContext.
4. AFAContext: A Novel Synthetic Dataset for Active Feature Acquisition
In this section, we introduce a novel dataset for AFA, called AFAContext. To motivate its design, we first describe the commonly used CUBE dataset (Rückstieß et al., 2013; Shim et al., 2018; Zannone et al., 2019), illustrated in Figure 0(a). CUBE consists of 20-dimensional real-valued vectors divided into 8 classes. For each class, three informative features are drawn from a Gaussian distribution $\mathcal{N}(\mu,0.3)$, with class-specific means $\mu$, while the remaining features are sampled from $\mathcal{N}(0.5,0.3)$ as noise. The informative features appear in different positions depending on the class, as shown by the colored cells in the figure (numbers correspond to the means). This dataset highlights the benefits of instance-specific (dynamic) feature selection, since different features are informative for different examples. However, it does not meaningfully differentiate between greedy and non-greedy methods: early in the selection process, methods must guess randomly among feature groups, and once a single informative feature is found, both greedy and non-greedy approaches can easily identify the rest. Thus, while CUBE supports evaluation of dynamic selection, it is less effective for evaluating the added value of lookahead or non-myopic strategies.
Figure 0(b) shows eight example instances from the AFAContext dataset. This dataset includes two one-hot encoded context features, $f_{1}$ and $f_{2}\in\{0,1\}$, which indicate which group of features is informative. When $f_{1}=1$, the ten features $f_{3},\dots,f_{12}$ are informative and follow the same structure as the first ten features of the CUBE dataset shown in Figure 0(a). The remaining ten features contain no information (random noise). In contrast, when $f_{2}=1$, the last ten features $f_{13},\dots,f_{22}$ are informative and also mirror the structure of CUBE. An optimal acquisition policy should therefore first select the context features $f_{1}$ and $f_{2}$ to identify the relevant context group, and then acquire the corresponding informative features. This dataset highlights both the value of instance-specific feature selection (as in CUBE) and the importance of non-greedy feature acquisition. This is because the context features $f_{1}$ and $f_{2}$ do not offer predictive value on their own and are therefore overlooked by greedy methods that focus solely on maximizing short-term performance.
5. Experiments
(a) AFAContext (Budget = 10)
(b) cube (Budget = 10)
(c) diabetes (Budget = 15)
(d) FashionMNIST (Budget = 30)
(e) physionet (Budget = 15)
(f) MNIST (Budget = 30)
(g) miniboone (Budget = 15)
Figure 2. Results for all methods across the seven datasets using the shared classifier. The reported results correspond to the large budget setting. See Appendix A.4 for other budgets and settings.
In this section, we present the experimental evaluation of the methods described in Section 3.1. Additional experimental results are illustrated in Appendix A.4. All implementation details, including model architectures and hyperparameters, are provided in Appendix A.
For each dataset, we consider three different budget values $b$ (small, medium and large) to illustrate the influence of this parameter. The choice of $b$ affects mainly the RL methods, as it is incorporated into the training process through the definition of the MDP. In addition, we conduct experiments using a shared classifier between all methods. This classifier is pretrained on the training set and then used by all methods during evaluation on the test set, ensuring a fair comparison. However, some methods include a built-in classifier that is trained jointly with the policy. To highlight the impact of this design choice, we also report results using the built-in classifier for the methods that support it.
In the main text, we present results for the large budget setting using the shared classifier (refer to the figures for the specific budget values used for each dataset). The results for all other settings are provided in the Appendix A.
@@ 5. Experiments
Finally, for static methods, we identify the $b$ most informative features in advance on the training data. During evaluation, CAE-S randomly selects features from this set, one at a time, until the budget is exhausted, since it does not provide a relative ranking of feature importance. In contrast, PT-S selects the features sequentially in order of the estimated importance. This setup allows for a fair comparison with dynamic methods, which naturally select features one at a time.
5.1. Results on Synthetic Data
Figures 1(a)1(b) present results for the large budget setting for the two synthetic datasets, AFAContext and CUBE, respectively. For CUBE, we observe that the discriminative greedy methods, GDFS-DG and DIME-DG, achieve the strongest performance. In particular, they are the only dynamic methods that outperform the static baseline PT-S, highlighting that static methods can be highly competitive. However, given the structure of this dataset (see Figure 0(a)), it is reasonable to expect dynamic methods to outperform static ones under lower noise levels.
For AFAContext, the RL-based methods ODIN-MFRL and OL-MFRL significantly outperform all others, demonstrating that reinforcement learning approaches are well suited to capturing the non-greedy structure of this dataset. In contrast, all greedy methods perform poorly as expected. Some RL methods also show worse performance, probably due to convergence challenges. The oracle-based non-greedy method AACO also performs poorly for this dataset. This is likely due to the heuristic proposed in the original paper (Valancius et al., 2024), which enables one-at-a-time feature selection but reduces the method’s non-greedy behavior. Finally, the model-free variant of ODIN (ODIN-MFRL) outperforms its model-based counterpart (ODIN-MBRL). This suggests that model-based RL may not offer clear advantages for the AFA setting, especially when a large amount of data is available. However, there may still be scenarios where model-based training proves beneficial, depending on the dynamics of the environment or the nature of the task.
5.2. Results on Real-World Data
Figures 1(c)1(g) present results for all real-world datasets in the large budget setting. We observe that the discriminative greedy methods, GDFS-DG and DIME-DG, consistently perform among the best across all datasets. In particular, they significantly outperform other methods on FashionMNIST and MNIST. In contrast, the greedy generative method EDDI-GG performs noticeably worse. Although it aims to estimate the same underlying quantity (CMI), its performance is likely to be hindered by the limitations of the generative PVAE model used by this method. Static methods also perform well in some cases. For example, PT-S ranks among the top performers on diabetes, miniboone, and PhysioNet, suggesting that selecting the same features for all instances is effective for these datasets. On the other hand, for datasets such as MNIST and FashionMNIST, dynamic feature selection, where different features may be chosen for each instance, is substantially more beneficial.
The non-greedy methods perform relatively poorly across most of the real-world datasets. Moreover, as discussed in the previous subsection, many of the non-greedy methods perform poorly on AFAContext, where they should in principle do well. This indicates two things: (i) many real-world datasets do not exhibit non-greedy structure (as our synthetic dataset AFAContext does by construction), (ii) non-greedy methods are more difficult to train due to the increased complexity of the methods, which is typically required because finding good non-greedy policies is significantly harder than greedy selection.
Finally, we observe that AACO performs particularly poorly on FashionMNIST and MNIST. This is likely due to the limited scalability of the method in high-dimensional settings, as it relies on approximating an enumeration over all possible feature subsets by random sampling (Valancius et al., 2024). As the dimensionality increases, the quality of this approximation deteriorates, leading to reduced performance.
5.3. Compute Time Comparison
Table 3 shows the time it took to train and evaluate each method. The training time also includes pretraining time, for those methods that have such a stage. The large standard deviations are due to different datasets.
Table 3. Compute time (in seconds) comparison across methods. Values are shown as mean ± standard deviation over multiple datasets, splits, and budgets.
Method Train Time (sec) Eval. Time (sec)
JAFA-MFRL $13523\pm 2769$ $68\pm 59$
ODIN-M(F&B)RL $9271\pm 6921$ $\mathbf{23}\pm 10$
OL-MFRL $2196\pm 2267$ $\mathbf{19}\pm 6$
EDDI-$G$ $7333\pm 5823$ $2032\pm 3495$
DIME-$G$ $1697\pm 1936$ $\mathbf{19}\pm 7$
GDFS-DG $3826\pm 4395$ $\mathbf{18}\pm 6$
PT-S $268\pm 296$ $\mathbf{19}\pm 7$
CAE-S $494\pm 473$ $19\pm 7$
AACO $\mathbf{10}\pm 3$ $15676\pm 7$
RL methods generally need more computational resources while training, especially JAFA-MFRL which trains an RNN. However, their output is a simple neural network that is very fast at evaluation time. In contrast, methods like AACO require almost no training time at all but are orders of magnitude slower at inference time.
6. Conclusion
In this work, we study the problem of Active Feature Acquisition (AFA), which aims to balance predictive accuracy with the cost of acquiring features. To allow a fair and comprehensive evaluation of the AFA methods, we introduce AFABench, the first standardized benchmark framework for this task. AFABench includes a diverse collection of synthetic and real-world datasets, supports a wide range of acquisition policies, and provides a modular design to promote extensibility. We implemented and evaluated representative algorithms from major AFA paradigms—including static, greedy, and reinforcement learning-based approaches—and introduced the novel synthetic dataset AFAContext to assess the lookahead capabilities of different methods. Our empirical analysis highlights key trade-offs among AFA strategies: (i) Although non-greedy methods are capable of learning non-myopic policies that can outperform greedy selection, this is not guaranteed, largely due to the increased complexity and potential instability of these approaches. Moreover, as our results show, many real-world datasets do not exhibit strong non-greedy structure. In such cases, greedy methods, particularly discriminative ones, offer a strong alternative with significantly shorter training time. (ii) Dynamic, instance-wise feature selection is often advantageous, especially for datasets with high variability across samples. However, for some datasets, static selection performs comparably well. In addition to their competitive performance, static methods are typically more efficient to train. We hope that this benchmark will serve as a basis for future research in cost-sensitive learning and more effective feature acquisition strategies.
Acknowledgments
Appendix A Experimental Details
A.1. Implementation Details
The shared classifier used to produce the results in Section 5 is a multilayer perceptron with the following configuration:
  • Dropout rate: $0.1$
@@ A.1. Implementation Details
  • Two hidden layers, each with $128$ units and ReLU activations.
  • Masking probability sampled from $\mathcal{U}(0.75,0.99)$ for (Fashion)MNIST
  • Masking probability sampled from $\mathcal{U}(0,0.9)$ for the remaining datasets
It was trained until the validation loss on samples with minimum masking (i.e., probability $0.25$ on (Fashion)MNIST and 0 on the other datasets) started increasing. Then, the checkpoint where the model had the lowest validation loss was saved.
@@
In general, hyperparameters are always set to the values found in the paper or original implementation. In cases where a hyperparameter was not listed publicly, ”standard” values have been chosen. Examples of this include setting the learning rate to $0.001$ and $\tau=0.005$ when soft-updating target networks.
A.1.1. Common across all methods
Masking probabilities Several methods need a pretraining stage where the features are masked with varying probabilities. In our pretraining runs, we always use a fixed masking probability per batch of data, and the probability itself is sampled from the distribution
@@ A.1.1. Common across all methods
Masking all features is not useful, and since many methods can achieve reasonable performance on only very small percentage of the total number of features in (Fashion)MNIST, we increase the masking probability.
Consistent loss function We use weighted cross-entropy for all methods, using class probabilities from the training set.
A.1.2. Common across RL methods
@@
Computations were performed on the GPU as much as possible. This means that both the agent (with replay buffers) and the environment were placed on the GPU.
Fair number of pretraining epochs All of the RL methods have a pretraining stage, where a model is trained using supervised learning. Pretraining is always performed until convergence, always saving the model with the lowest validation loss on features masked with the minimum masking probability. As listed above, for (Fashion)MNIST this is 25% and for the other datasets it is 10%.
No discounting We do not use discounting (therefore, setting $\gamma=1$) since the MDP is episodic and the episodes are short ($\leq 30$ steps).
Consistent batch size, number of agents, and training epochs During the training stage, data is collected in consistent batches and with the same number of parallel agents for all RL methods. “Agents” in this case do not refer to separate networks, but simply to the fact that experience is collected in parallel rollouts. We use the same values as in (Shim et al., 2018), i.e.,
  • batch size $512$, with $128$ parallel agents, for all datasets except (Fashion)MNIST. This means that each agent receives a batch of size $4$.
  • batch size $64$ and $2$ parallel agents for (Fashion)MNIST. This modification is done to allow everything to fit on the GPU.
The number of batches that each agent receives is 10000 for all datasets, except for (Fashion)MNIST which needed more batches (30000).
A.1.3. JAFA-MFRL
@@
  • Original JAFA specifies the hidden layers for the reading block cells of the set encoder introduced in (Vinyals et al., 2016) as $[32,32]$, but not the hidden layers of the writing block. We choose to set this to $[32,32]$ as well.
  • We use $5$ processing steps for the RNN.
  • Original JAFA uses a constant masking probability of $0.5$, while ours varies.
  • Original JAFA uses hidden layers $[32,32]$ for the classifier. We do this for all datasets, except for (Fashion)MNIST where we found that the classifier severely underfitted. In this case, we increased the model complexity to $[128,128]$ instead.
@@
  • Original JAFA uses 4-step Q-learning. Since TorchRL does not make this readily available, we instead use a TD($\lambda$) value estimator with $\lambda=0.75$. This value is chosen since it gives the largest weight to the 4-step return (see Appendix A.3).
@@
  • This is by far the slowest of the RL methods, due to the RNN.
@@ A.1.4. ODIN-MBRL
  • Original ODIN masks features with probability sampled from $\mathcal{U}(0,1)$, we have different probabilities.
  • The scaling factor for the KL loss term when training the PVAE is set to $0.1$ for all datasets, except for AFAContext where it is set to $0.01$. We tried the default value of $1$ first but found that it led to posterior collapse.
@@
  • Original ODIN does not describe the number of parallel agents and the batch size that they use. We use consistent values across all RL methods.
  • In RL, it is common to perform multiple gradient steps on the same batch of data. We do a single pass, which is also what the original JAFA does.
  • For PPO we use an entropy bonus coefficient of $0.01$, the default value in TorchRL.
Original ODIN uses single linear layers for both the value network and the policy network of the PPO agent, receiving a latent vector from the PVAE as input. We found that this can cause underfitting in some cases, which is why we changed the networks to hidden layers with dimensionality $[32,32]$ for comparisons on the AFAContext dataset.
A.1.5. ODIN-MFRL
ODIN-MFRL was trained exactly the same way as ODIN-MBRL, expect that it disables generative rollouts.
A.1.6. OP-MFRL
In the original implementation, there is no pretraining stage for training the classifier separately. However, the reward that the agent receives is only as useful as the classifier it is based on, and we found that a majority of the training time was ”wasted” on waiting for the classifier to improve. Hence, we added a pretraining stage, improving training efficiency.

### v3

@@ D.4. Static (non-adaptive) feature selection methods
Static methods are restricted to the hard-budget setting. The reason is that they do not perform instance-wise selection, which makes the soft-budget setting irrelevant.
@@
To adapt static feature selection methods to the active feature acquisition setting considered in this paper, we first identify a fixed set of the most informative features in advance on the training data. During evaluation, static methods use the same feature subset for all test instances under a hard budget $b$. For CAE-S, since it does not provide a relative ranking of feature importance, we use the CAE selector only to identify the optimal feature subset for the maximum budget, and during evaluation, features from this subset are selected sequentially in a fixed index-based order until the budget is exhausted. In contrast, PT-S selects the features sequentially in order of the estimated importance. This setup allows for a fair comparison with dynamic methods, which naturally select features one at a time.
Appendix E Experimental Details
E.1. Implementation Details
E.1.1. External classifiers
Two different external classifiers were trained. They both followed a standard supervised learning loop, with the exception of using masking on each batch retrieved from the dataset. For each batch, every feature was masked uniformly with some probability that was drawn from a uniform distribution. The range of this distribution depended on the classifier and dataset. Checkpointing was used to save the model with lowest validation loss on batches with minimum masking probability.
The first classifier, which was used on all datasets except Imagenette, was a multilayer perceptron with the following common properties shared across datasets:
@@ E.1.1. External classifiers
  • Two hidden layers, each with $128$ units and ReLU activations.
Other properties, like batch size, number of training epochs, and masking probability, were slightly tuned for each dataset, to ensure good classification performance.
The second classifier, which was used on Imagenette, was a Vision Transformer with the following configuration:
  • Backbone: pretrained ViT (vit_small_patch16_224)
  • Input resolution: $224\times 224$
  • Patch size: $16$
@@
  • Optimizer: Adam
  • Training epochs: $200$
  • Masking probability sampled from $\mathcal{U}(0,0.9)$
  • Learning rate scheduling: ReduceLROnPlateau applied on the validation loss with a reduction factor of $0.2$, patience of $3$, initial learning rate of $1\times 10^{-5}$, and minimum learning rate of $1\times 10^{-8}$
E.1.2. Method implementations
@@
In general, hyperparameters were always set to the values found in the paper or original implementation. In cases where a hyperparameter was not listed publicly, ”standard” values have been chosen. Examples of this include setting the learning rate to $0.001$ and $\tau=0.005$ when soft-updating target networks for RL.
E.1.3. Common across all methods
Consistent masking probabilities during pretraining
Several methods needed a pretraining stage where the features are masked with varying probabilities. In our pretraining runs, we always used a fixed masking probability per batch of data, and the probability itself was sampled from the distribution
@@ E.1.3. Common across all methods
Masking all features is not useful, and since many methods can achieve reasonable performance on only very small percentage of the total number of features in (Fashion)MNIST, we increased the masking probability.
Consistent loss function For methods that required a loss function, we used weighted cross-entropy, using class probabilities from the training set.
E.1.4. Common across RL methods
@@
Computations were performed on the GPU as much as possible. This means that both the agent (with replay buffers) and the environment were placed on the GPU. One exception to this was OL-MFRL which required the replay buffer to be placed in RAM when training on (Fashion)MNIST.
Pretraining until convergence All of the RL methods have a pretraining stage, where a model is trained using supervised learning. Pretraining was always performed until convergence by using early stopping, always saving the model with the lowest validation loss on features masked with the minimum masking probability. As listed above, for (Fashion)MNIST this is 25% and for the other datasets it is 10%.
No discounting We did not use discounting (therefore setting $\gamma=1$) since the MDP is episodic.
Consistent batch size, number of agents, and training epochs During the training stage, data was collected in consistent batches and with the same number of parallel agents for all RL methods. “Agents” in this case does not refer to separate networks, but simply to the fact that experience was collected in parallel rollouts. We used the same values as in (Shim et al., 2018), i.e.,
  • batch size $512$, with $128$ parallel agents, for all datasets. This means that each agent received a batch of size $4$.
Furthermore, the total number of batches used was always 10000. Although some methods would have possibly benefited from more training, we wanted to avoid excessive hyperparameter tuning. Instead, by giving the exact same amount of data to each method, sample efficiency became a part of the evaluation.
E.1.5. JAFA-MFRL
@@
  • Original JAFA specified the hidden layers for the reading block cells of the set encoder introduced in (Vinyals et al., 2016) as $[32,32]$, but not the hidden layers of the writing block. We chose to set this to $[32,32]$ as well.
  • We used $5$ processing steps for the RNN.
  • Original JAFA used a constant masking probability of $0.5$, while ours varied.
  • Original JAFA used hidden layers $[32,32]$ for the classifier. We used this for all datasets, except for (Fashion)MNIST where we found that the classifier severely underfitted. In this case, we increased the model complexity to $[128,128]$ instead.
@@
  • Original JAFA used 4-step Q-learning. Since TorchRL does not make this readily available, we instead used a TD($\lambda$) value estimator with $\lambda=0.75$. This value was chosen since it gives the largest weight to the 4-step return (see Appendix E.3).
@@
  • This is by far the slowest of the RL methods, due to the RNN. See Section˜A.4.
@@ E.1.6. ODIN-MBRL
  • Original ODIN masked features with probability sampled from $\mathcal{U}(0,1)$, we have different probabilities.
  • The scaling factor for the KL loss term when training the PVAE was set to $0.1$ for all datasets. We tried the default value of $1$ first but found that it led to posterior collapse.
  • Similarly to JAFA, hidden layers $[128,128]$ were used for (Fashion)MNIST and $[32,32]$ for the remaining datasets.
@@
  • Original ODIN does not describe the number of parallel agents and the batch size that they use. We used consistent values across all RL methods.
  • For PPO we used an entropy bonus coefficient of $0.01$, the default value in TorchRL.
E.1.7. ODIN-MFRL
ODIN-MFRL was trained in exactly the same way as ODIN-MBRL, expect that it disabled generative rollouts.
E.1.8. OL-MFRL+no-mask
In the original implementation, there was no pretraining stage for training the classifier separately. However, the reward that the agent receives is only as useful as the classifier it is based on, and we found that a majority of the training time was ”wasted” on waiting for the classifier to improve. Hence, we added a pretraining stage, improving training efficiency.